(10%) Problem 7: NASA uses a device often called the "vomit comet" to simulate the effects of low gravity in space during astronaut training. The device moves in circular arcs at a speed of v = 213 m/s.



What is the radius the pilot must fly to create weightlessness as they go around the top of the circular arc? Give your answer in meters.

1
Expert's answer
2022-01-21T08:25:22-0500

To reach weightlessness, the centrifugal acceleration must be equal to the acceleration due to gravity:


"a=g=\dfrac {v^2}R→R=\dfrac {v^2}g=4629\text{ m}."